# Leveraging Sub-Optimal Data for Human-in-the-Loop Reinforcement Learning

### v1

@@ Abstract
To create useful reinforcement learning (RL) agents, step zero is to design a suitable reward function that captures the nuances of the task. However, reward engineering can be a difficult and time-consuming process. Instead, human-in-the-loop (HitL) RL allows agents to learn reward functions from human feedback. Despite recent successes, many of the HitL RL methods still require numerous human interactions to learn successful reward functions. To improve the feedback efficiency of HitL RL methods (i.e., require less feedback), this paper introduces Sub-optimal Data Pre-training, SDP, an approach that leverages reward-free, sub-optimal data to improve scalar- and preference-based HitL RL algorithms. In SDP, we start by pseudo-labeling all low-quality data with rewards of zero. Through this process, we obtain “free” reward labels to pre-train our reward model. This pre-training phase provides the reward model a head start in learning, whereby it can identify that low-quality transitions should have a low reward, all without any actual feedback. Through extensive experiments with a simulated teacher, we demonstrate that SDP can significantly improve or achieve competitive performance with state-of-the-art (SOTA) HitL RL algorithms across nine robotic manipulation and locomotion tasks.
@@ 1 Introduction
In reinforcement learning (RL), an agent’s goal is to interact with an environment to maximize its total reward Sutton and Barto (2018). It is assumed that the environment provides an agent with a well-defined reward function that captures all task complexities. But where does this reward function actually come from? Reward functions are hand-engineered by humans in what often can be a tedious and non-trivial pursuit Booth et al. (2023). As the complexity of tasks increases, so does the time and effort required to design a suitable reward function. Further, there have been notable examples of reward misspecification, in which RL agents discovered and exploited unintended shortcuts in the reward function  Skalse et al. (2022). One notorious example is the CoastRunners game, in which the objective should be to finish a boat race as fast as possible. Instead, an RL agent gained the most reward by spinning its boat in a circle despite concurrently catching on fire and crashing into other boats Clark and Amodei (2016).
@@
A promising alternative is to learn reward functions directly from human feedback. In this paradigm, humans can provide feedback in the form of preferences or scalar signals, which can then used to learn a reward function that is consistent with human desires Daniel et al. (2014); Christiano et al. (2017); Lee et al. (2021); White et al. (2023). Despite recent progress, existing preference- and scalar-based HitL RL methods still suffer from high human labeling costs that can require thousands of human queries to learn an adequate reward function Christiano et al. (2017). This may invalidate the original goals of human-in-the-loop for reward design. Prior work attempts to mitigate this issue through several mechanisms, including active learning Settles (2009); Lee et al. (2021), data augmentation Shorten and Khoshgoftaar (2019); Park et al. (2022), semi-supervised learning Zhu (2005); Park et al. (2022), and meta-learning Muslimani et al. (2022); Hejna III and Sadigh (2023).
Alternatively, this work takes inspiration from data sharing in offline RL, in which the goal is to leverage prior data to improve performance on a given target task. In settings where there is an abundance of unlabeled (i.e., reward-free), low-quality data, one way to use this data is to simply label all such data with a reward of zero (i.e., the minimum reward for the task) Yu et al. (2021). This simple approach achieved promising results on a variety of robotic tasks in the offline RL setting. As mediocre or low-quality data is arguably the easiest type of data to obtain, and prior work has demonstrated how to leverage it in offline RL, this paper asks the question:
Can we leverage abundant sub-optimal, unlabeled data to improve learning in HitL RL methods?
@@
To that end, we present Sub-optimal Data Pre-training, SDP, a tool for HitL RL algorithms to increase human feedback efficiency. SDP leverages sub-optimal target task data by pseudo-labeling all transitions with the minimum reward for the task, which we assume to be zero (without loss of generality). The now pseudo-labeled sub-optimal data is used in two ways. First, we pre-train a regression-based reward model by applying standard supervised learning to minimize the mean squared loss. Intuitively, the pre-training component provides a “free” head start to the reward model, whereby the reward model can bias these low-quality transitions to have a lower reward value than other transitions. Second, we initialize the RL agent’s replay buffer with the sub-optimal data and make learning updates to the RL agent. This process changes the RL agent’s policy and provides new behaviors for the human to provide feedback on, relative to learning with no initial sub-optimal data. This ensures that when it is time for the human teacher to provide feedback, they will not be providing redundant feedback to the existing sub-optimal data. Afterward, we follow standard preference- or scalar-based HitL RL.
@@
This paper’s core contribution is showing that we can harness the availability of low-quality, reward-free data for HitL RL approaches by pseudo-labeling it with zero rewards and treating it as a free bias for learning reward models. Extensive experiments combining SDP with both scalar- and preference-based reward learning RL algorithms show that the addition of SDP can significantly improve feedback efficiency across complex tasks from both DeepMind Control (DMControl) Tassa et al. (2018) and Meta-world Yu et al. (2020) suites. We additionally show that SDP is not limited to leveraging sub-optimal data from the target task — sub-optimal data from different tasks can also be useful, highlighting the generality of SDP. Overall, this work takes an important step toward considering how HitL RL approaches can take advantage of readily-available sub-optimal data.
@@ 2 Related Work
This section highlights related work in HitL RL and in leveraging sub-optimal data.
2.1 Human-in-the-Loop RL
@@
Several approaches in HitL RL allow agents to leverage human feedback to adapt or learn new behavior. Learning from demonstration (LfD) Argall et al. (2009) is one such methodology that allows a human to provide examples of desired agent behavior. Human demonstration data has been used to shape the environment’s reward function Brys et al. (2015), develop a reward function from scratch Abbeel and Ng (2004), and bias the agent’s policy towards certain actions Taylor et al. (2011). Although demonstrations can be a rich source of feedback, they are often expensive to obtain and may require domain experts Dragan and Srinivasa (2012).
@@
Another approach is learning from preference-based feedback where a teacher provides preferences between two or more sets of agent behavior Christiano et al. (2017). Preference learning has been popularized in recent years as it can require less effort and expertise compared to providing demonstrations. To further reduce the amount of human interaction required, several strategies have been introduced. Recent work has combined preferences with demonstrations Ibarz et al. (2018); Bıyık et al. (2022), used unsupervised pre-training for policy initialization Lee et al. (2021), integrated semi-supervised learning and data augmentation techniques Park et al. (2022), applied uncertainty-based exploration strategies Liang et al. (2022), and leveraged labeled data from multiple tasks via a meta-learning approach Hejna III and Sadigh (2023). Despite its popularity, some argue that comparison feedback might not capture the full intricacies of human preferences, as oftentimes the human is limited to choosing between two options Daniel et al. (2014); White et al. (2023).
@@
As a result, another body of work focuses on learning from scalar feedback where human teachers can provide scalar signals to evaluate an agent’s behavior Knox and Stone (2009); Griffith et al. (2013); Loftin et al. (2016); MacGlashan et al. (2017); Warnell et al. (2018); White et al. (2023). Several works use scalar feedback to learn a reward model via regression Daniel et al. (2014); Cui and Niekum (2018); Cabi et al. (2020). Similarly, other works learn a human reinforcement function via regression Knox and Stone (2009, 2013); Warnell et al. (2018). In this setting, a human is assumed to provide scalar feedback that is representative of a behavior’s long-term value. A human reinforcement function is distinct from a reward function, and more closely resembles an action-value function in RL.
Despite the growing advances in HitL RL, there is no work, to the best of our knowledge, in understanding how reward-free, sub-optimal data can be leveraged to improve the feedback efficiency of preference- and scalar-based approaches.
2.2 Learning from Sub-Optimal Data
@@
SDP aims to leverage sub-optimal data for scalar- and preference-based HitL RL algorithms. However, learning from low-quality data or negative examples has been applied in other areas of reinforcement learning and imitation learning Chen et al. (2021); Tangkaratt et al. (2021). In standard RL, several works use sub-optimal demonstrations to initialize a policy Taylor et al. (2011); Hester et al. (2018); Gao et al. (2019). In offline RL, some approaches leverage sub-optimal transitions from multiple tasks and either assign reward labels according to the reward function of the target task Singh et al. (2020) or simply label them the minimum possible environmental reward Yu et al. (2022). In goal-conditioned RL, Hindsight-Experience-Replay (HER) Andrychowicz et al. (2017) uses failed episodes by treating them as a success with respect to a different goal. In inverse reinforcement learning (IRL) Arora and Doshi (2021), a constrained optimization formulation has been proposed that can accommodate both successful and failed demonstrations Shiarlis et al. (2016). Trajectory-ranked Reward Extrapolation (T-REX) Brown et al. (2019) can leverage sub-optimal demonstrations in IRL, making use of ranked demonstrations to learn a reward function. Later work improves upon T-REX by learning a policy via behavioral cloning on the demonstrations. They then automatically generate ranked trajectories by adding increasing amounts of noise to the learned policy Brown et al. (2020).
@@ 3 Background
In the RL paradigm, agents interact with an environment with the goal of maximizing a scalar reward signal. This interaction process is modeled as a Markov Decision Process (MDP) which consists of $\langle\mathcal{S},\mathcal{A},T,r,\gamma\rangle$. At every time-step $t$, the agent receives a state $s_{t}\in\mathcal{S}$ from the environment and chooses an action $a_{t}\in\mathcal{A}$. The environmental transition function, $T$, determines the probability of transitioning to state $s_{t+1}$ and receiving reward $r_{t+1}$, given the agent was in state $s_{t}$ and executed action $a_{t}$. The environment then provides the agent with this scalar reward $r_{t+1}$. The agent attempts to learn a policy, $\pi:\mathcal{S}\rightarrow\mathcal{A}$, that maximizes the return $G=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$, which is defined as the expected sum of discounted future rewards with discount factor $\gamma\in$ $[0,1)$.
@@ 3.1 Reward Learning from Human Feedback
This paper assumes that we are in a reward-free paradigm, an MDP/R setting, where our goal is to learn a good policy while also learning a reward function from human feedback. We follow the standard reward learning framework that uses supervised learning to learn such a reward function, $\hat{r}_{\theta}$ Daniel et al. (2014); Christiano et al. (2017). In both scalar- and preference-based feedback settings, we consider segments of trajectories $\sigma$. Each segment consists of a sequence of states and actions {$s_{t},a_{t},s_{t+1},a_{t+1},...,s_{t+k},a_{t+k}$}.
@@
In preference-based learning, two segments, $\sigma^{0}$ and $\sigma^{1}$, are compared by a teacher, yielding $y\in$ $0,.5,1$}. Specifically, if the teacher preferred segment $\sigma^{1}$ over segment $\sigma^{0}$, then $y$ is set to 1, and if the converse is true $y$ is set to 0. If both segments are equally preferred, then $y$ is set to .5.5.5.5. As feedback is collected, it is stored as tuples ($\sigma^{0},\sigma^{1},y$) in the reward model (RM) data set $D_{RM}$. Then, we follow the Bradley-Terry model Bradley and Terry (1952) to define a preference predictor using the reward function $\hat{r}_{\theta}$:
@@
In general, if $\sigma^{i}>\sigma^{j}$ in equation 1, then the segment $\sigma^{i}$ is preferred by the teacher over segment $\sigma^{j}$. Intuitively, this model assumes that the probability of the teacher preferring a segment depends exponentially on the total sum of predicted rewards along the segment. To train the reward function, we can use supervised learning where the teacher provides the labels $y$. More specifically, we update $\hat{r}_{\theta}$ by minimizing the standard binary cross-entropy objective:
$L^{CE}(\theta,D)=-\ E_{(\sigma^{0},\sigma^{1},y)\sim D}\bigl{[}(1-y)\>\text{log}P_{\theta}(\sigma^{0}>\sigma^{1})\\ +\;y\>\text{log}P_{\theta}(\sigma^{1}>\sigma^{0})\bigr{]}$ (2)
@@
The primary difference between scalar and preference-based reward learning is that in scalar-based learning, the human teacher assigns numerical ratings to segments of trajectories one at a time. In this setting, the comparisons between segments are implicit. More concretely, a teacher assigns a scalar value $y$ to a segment $\sigma^{i}$, and as feedback is collected, it is stored as tuples ($\sigma^{i},y$) in the reward model data set $D_{RM}$. We then apply standard regression and update $\hat{r}_{\theta}$ by minimizing the mean squared error:
$L^{MSE}(\theta,D)=E_{(\sigma^{i},y)\sim D}\bigl{[}(y-\hat{r}_{\theta}(\sigma^{i}))^{2}\bigr{]}$ (3)
@@ 4 Sub-optimal Data Pre-training
In this section, we present SDP, a tool to improve the feedback efficiency for HitL RL. We take the approach of pseudo-labeling all transitions from a set of sub-optimal demonstrations with the lowest possible environmental reward (we assume to be zero). The goal of SDP is then to use this pseudo-labeled data to create a pessimistic prior for rewards models in HitL RL methods (see Figure 1). Its simplicity enables SDP to be used in conjunction with any off-the-shelf HitL RL algorithm that learns a reward function from feedback. Algorithm 1 contains the complete pseudocode.
4.1 Pessimistic Prior for Reward Learning
@@
SDP comprises two phases: (1) the reward model pre-training phase and (2) the agent update phase. In the reward model pre-training phase, we first gather a data set, $D_{\text{sub}}$, of $N$ sub-optimal state, action transitions. We then pseudo-label all transitions in $D_{\text{sub}}$ with rewards of 00, resulting in $D_{\text{sub}}=\{s_{i},a_{i},0\}^{N}_{i=1}$. $D_{\text{sub}}$ is then used to optimize the reward model $\hat{r}_{\theta}$ with the mean squared loss in Equation 3 (see lines 2-4 in Algorithm 1). As a result, the reward model $\hat{r}_{\theta}$ becomes pessimistic because it learns to associate all sub-optimal transitions with a low reward. Without such a prior, the reward model initially has random estimates for the sub-optimal transitions. The only way to improve such estimates is to obtain teacher feedback. Therefore, by pseudo-labeling the sub-optimal transitions with 00, we obtain a helpful reward bias before receiving any teacher feedback.
@@
Next, in the agent update phase, we initialize the RL agent’s replay buffer $D_{\text{agent}}$ with $D_{\text{sub}}$ (see line 6 in Algorithm 1). The RL agent then briefly interacts with its environment and performs gradient updates according to its loss functions (see lines 7-12 in Algorithm 1). The agent update process changes the RL agent’s policy and generates new transitions, which are then stored in both the agent’s replay buffer $D_{\text{agent}}$ (see line 9 in Algorithm 1) and the reward model’s data set $D_{\text{RM}}$ (see line 10 in Algorithm 1). It is important to note that in standard scalar- and preference-based reward learning, we query the teacher for feedback on trajectory segments sampled from $D_{\text{RM}}$. Therefore, adding new transitions into $D_{\text{RM}}$ during the agent update phase is necessary to ensure that the teacher does not provide redundant feedback to the original sub-optimal transitions (as $D_{\text{RM}}$ was empty prior to the agent update phase). When it is time for the teacher to provide their first set of feedback (see line 13 in Algorithm 1), the feedback can cover a different region of the state and action space, relative to the original sub-optimal data. We empirically show that the agent update phase changes the RL agent’s policy by performing policy rollouts and analyzing the differences in state distributions (see Figure 5 in Appendix B).
@@
One may immediately think to ask: is labeling sub-optimal transitions with an incorrect reward problematic? Using incorrect reward labels does induce some incorrect bias for both the reward model and the RL agent’s value network. However, as the transitions are sub-optimal, the bias for using an incorrect reward is low and does not greatly influence learning. Figure 6 in Appendix B shows that the majority of the reward values for sub-optimal transitions gathered through a random policy do have true rewards close to 0. In addition, by using the sub-optimal transitions, we increase the overall amount of data used by both models, which can decrease the models’ variance Yu et al. (2022).
@@
Figure 1: Overview of SDP: After obtaining sub-optimal target task data, we pseudo-label this data with rewards of zero. We then pre-train our reward model $\hat{r}_{\theta}$ using this data set. During the agent update phase, we initialize our RL agent’s replay buffer with the same pseudo-labeled data set. We then interact in the environment and make learning updates to obtain new behaviors for a teacher to give feedback.
Figure 2: Scalar feedback experiments. In all DMControl experiments, SDP significantly outperforms R-PEBBLE and Deep TAMER
(p $<$ 0.05) and achieves comparable performance to the SAC oracle.
Figure 3: Preference feedback experiments in DMControl and Metaworld suites. The bar plots show AUC +/- 95% confidence intervals.
* indicates that SDP + the base preference learning algorithm achieves a statistically greater score (p $<$ 0.05) than the base preference learning algorithm alone (i.e., without SDP).
@@
Require: Reward model $\hat{r}_{\theta}\leftarrow\theta$ randomly initialized, Reward model data set $D_{\text{RM}}\leftarrow\emptyset$, RL agent with replay buffer $D_{\text{agent}}\leftarrow\emptyset$, Sub-optimal data set $D_{\text{sub}}$ with reward labels 00
@@
3:           Optimize $\hat{r}_{\theta}$ on $D_{\text{sub}}$ with $L^{MSE}$ (equ. 3)
@@
13:  Begin scalar- or preference-based reward learning using pre-trained $\hat{r}_{\theta}$, RL agent, and $D_{\text{agent}}$, $D_{\text{RM}}$
4.2 Implementation Details of SDP
For ease of obtaining sub-optimal data for SDP, we used state, action transitions from a random policy. We further note that we do not require explicit access to a sub-optimal policy; we only require state, action transitions from said policy. For all the experiments in Section 5.2, we used 50,000 transitions. Moreover, all reward model hyperparameters remained the same during both SDP phases (i.e., the reward model pre-train and agent update phases) as well as during the standard scalar- or preference-based reward learning that followed.

### v2

@@ Abstract
To create useful reinforcement learning (RL) agents, step zero is to design a suitable reward function that captures the nuances of the task. However, reward engineering can be a difficult and time-consuming process. Instead, human-in-the-loop RL methods hold the promise of learning reward functions from human feedback. Despite recent successes, many of the human-in-the-loop RL methods still require numerous human interactions to learn successful reward functions. To improve the feedback efficiency of human-in-the-loop RL methods (i.e., require less human interaction), this paper introduces Sub-optimal Data Pre-training, SDP, an approach that leverages reward-free, sub-optimal data to improve scalar- and preference-based RL algorithms. In SDP, we start by pseudo-labeling all low-quality data with the minimum environment reward. Through this process, we obtain reward labels to pre-train our reward model without requiring human labeling or preferences. This pre-training phase provides the reward model a head start in learning, enabling it to recognize that low-quality transitions should be assigned low rewards. Through extensive experiments with both simulated and human teachers, we find that SDP can at least meet, but often significantly improve, state of the art human-in-the-loop RL performance across a variety of simulated robotic tasks.
@@ 1 Introduction
In reinforcement learning (RL), an agent’s objective is to interact with an environment and maximize its total (discounted) expected reward. The reward hypothesis further maintains that a well-specified reward function is sufficient for an agent to learn to solve a task (Sutton & Barto, 2018). However, defining a reward function that precisely captures all task complexities is often tedious and non-trivial (Booth et al., 2023). There have been notable examples of reward misspecification, in which RL agents discovered and exploited unintended shortcuts in the reward function (Skalse et al., 2022). One notorious example is the CoastRunners game, in which the goal should be to finish a boat race as fast as possible — an RL agent instead gained the most reward by spinning its boat in a circle despite concurrently catching on fire and crashing into other boats (Clark & Amodei, 2016).
@@
A promising alternative is to learn reward functions directly from human feedback. In this paradigm, humans can provide feedback in the form of preferences or scalar signals, which can then be used to learn a reward function that is consistent with human desires (Daniel et al., 2014; Christiano et al., 2017). Despite recent progress, existing preference- and scalar-based RL methods still suffer from high human labeling costs that can require thousands of human queries to learn an adequate reward function (Christiano et al., 2017). Prior work attempts to mitigate this issue through several mechanisms, including active learning (Lee et al., 2021a), data augmentation (Park et al., 2022), semi-supervised learning (Park et al., 2022), and meta-learning (Hejna III & Sadigh, 2023).
Alternatively, our work draws on recent advances in offline RL that have demonstrated the value of low-quality data (Yu et al., 2021). However, its potential in human-in-the-loop RL remains unexplored. As low-quality data is often readily accessible or easy to obtain, this work addresses this gap by asking the question:
Can we leverage sub-optimal, unlabeled data
to improve learning in human-in-the-loop RL methods?
@@
To that end, we present Sub-optimal Data Pre-training, SDP, a tool for human-in-the-loop RL algorithms to increase human feedback efficiency. SDP leverages sub-optimal trajectories by pseudo-labeling all transitions with the minimum environment reward. The now pseudo-labeled sub-optimal data serves two purposes. First, we pre-train a regression-based reward model by applying standard supervised learning to minimize the mean squared loss. Intuitively, this pre-training provides the reward model a head start, biasing it towards assigning lower reward values to these low-quality transitions. Second, we initialize the RL agent’s replay buffer with the sub-optimal data and make learning updates to the RL agent. This process changes the RL agent’s policy and provides different behaviors for the human to provide feedback on (relative to learning with no initial sub-optimal data). This ensures that when the human teacher provides feedback, their time is used efficiently, avoiding redundant feedback on the existing sub-optimal data. Afterward, we follow the standard preference- or scalar-based RL protocol.
@@
This paper’s core contribution is showing that we can harness the availability of low-quality, reward-free data for human-in-the-loop RL approaches by pseudo-labeling it with minimum rewards and treating it as a prior for learning reward models. We first validate the utility of SDP in extensive simulated teacher experiments, combining it with four scalar- and preference-based RL algorithms. These experiments show that SDP significantly improves the feedback efficiency in complex tasks from both the DeepMind Control (DMControl) (Tassa et al., 2018) and Meta-World (Yu et al., 2020) suites. Crucially, we further highlight the real-world applicability of SDP by demonstrating its success with human teachers in a $16$-person user study. Overall, this work takes an important step toward considering how human-in-the-loop RL approaches can take advantage of readily available sub-optimal data.
@@ 2 Related Work
Human-in-the-Loop RL
@@
Several approaches in human-in-the-loop RL allow agents to leverage human feedback to adapt or learn new behavior. Learning from demonstration is one such methodology that allows a human to provide examples of desired agent behavior (Argall et al., 2009). Human demonstration data has been used to shape the environment’s reward function (Brys et al., 2015), develop a reward function from scratch (Abbeel & Ng, 2004), or bias the agent’s policy towards certain actions (Taylor et al., 2011). Although demonstrations can be a rich source of feedback, they are often expensive to obtain and may require domain experts (Dragan & Srinivasa, 2012).
@@
Another approach is learning from preference-based feedback where a teacher provides preferences between two or more sets of agent behavior (Christiano et al., 2017). Preference learning has been popularized in recent years as it can require less effort and expertise compared to providing demonstrations. To further reduce the amount of human interaction required, several strategies have been introduced. This has included combining preferences with demonstrations (Ibarz et al., 2018; Bıyık et al., 2022), unsupervised pre-training (Lee et al., 2021a), bi-level optimization (Liu et al., 2022), semi-supervised learning (Park et al., 2022), data augmentation (Park et al., 2022), uncertainty-based exploration (Liang et al., 2022), meta-learning (Hejna III & Sadigh, 2023), and active learning approaches (Hu et al., 2024). Despite its popularity, some argue that comparison feedback might not capture the full intricacies of human preferences, as oftentimes the human is limited to choosing between two options (Daniel et al., 2014; White et al., 2024).
@@
As a result, another body of work focuses on learning from scalar feedback where human teachers can provide scalar signals to evaluate an agent’s behavior (Knox & Stone, 2009; Griffith et al., 2013; Loftin et al., 2016; White et al., 2024). Several works use scalar feedback to either learn a reward model (Daniel et al., 2014; Cabi et al., 2020) or an action-value function (Knox & Stone, 2009; 2013; Warnell et al., 2018) via regression.
Learning from Sub-Optimal Data
@@
SDP aims to leverage sub-optimal data for scalar- and preference-based RL algorithms. However, learning from low-quality data or negative examples has been applied in other areas of RL and imitation learning (Chen et al., 2021; Tangkaratt et al., 2021). In standard RL, several works use sub-optimal demonstrations to initialize a policy (Taylor et al., 2011; Hester et al., 2018; Gao et al., 2019). In goal-conditioned RL, Hindsight-Experience-Replay uses failed episodes by treating them as a success with respect to a different goal (Andrychowicz et al., 2017). In inverse reinforcement learning (IRL), Shiarlis et al. (2016) proposed a constrained optimization formulation that can accommodate both successful and failed demonstrations. Brown et al. (2019) makes use of ranked demonstrations to learn a reward function in IRL. Later work in IRL automatically generates ranked trajectories by adding increasing amounts of noise to a learned policy (Brown et al., 2020). Lastly, in offline RL, Singh et al. (2020) leverages sub-optimal transitions from multiple prior tasks and assigns reward labels according to the current task reward function. Our work is most closely related to Yu et al. (2022), which leverages reward-free, sub-optimal data in the offline RL setting by pseudo-labeling all transitions with zero and adding them to the RL agent’s replay buffer. However, we found that directly applying this approach to the human-in-the-loop RL setting was ineffective (see Appendix D.5).
@@ 3 Background
In the RL paradigm, agents interact with an environment to maximize the total (discounted) expected reward it can achieve. This interaction is modeled as a Markov Decision Process (MDP) which consists of $\langle\mathcal{S},\mathcal{A},T,r,\gamma\rangle$. At every time-step $t$, the agent receives a state $s_{t}\in\mathcal{S}$ from the environment and chooses an action $a_{t}\in\mathcal{A}$. The transition function, $T$, determines the probability of transitioning to state $s_{t+1}$ and receiving reward $r_{t+1}$, given the agent was in state $s_{t}$ and executed action $a_{t}$. The environment then provides the agent $r_{t+1}$. The agent attempts to learn a policy, $\pi:\mathcal{S}\rightarrow\mathcal{A}$, that maximizes the expected return $\mathbb{E}[G]=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$, which is defined as the expected sum of discounted future rewards with discount factor $\gamma\in$ $[0,1)$.
@@ 3.1 Reward Learning from Human Feedback
This paper assumes that we are in a reward-free paradigm, an MDP/R setting, where our goal is to (1) learn a reward function, $\hat{r}$, from human feedback and (2) learn a policy that maximizes the total (discounted) expected $\hat{r}$. We follow the standard reward learning framework that uses supervised learning to learn a parameterized reward function, $\hat{r}_{\theta}$, with parameters $\theta$ (Christiano et al., 2017). In both scalar- and preference-based settings, we consider trajectory segments $\sigma$, where $\sigma$ consists of a sequence of states and actions: {$s_{t},a_{t},s_{t+1},a_{t+1},...,s_{t+k},a_{t+k}$}, with $k$ as the segment size.
@@
In preference-based learning, two segments, $\sigma^{0}$ and $\sigma^{1}$, are compared by a teacher, yielding $y\in$$0,0.5,1$}. Specifically, if the teacher preferred segment $\sigma^{1}$ over segment $\sigma^{0}$, then $y$ is set to 1, and if the converse is true $y$ is set to 0. If both segments are equally preferred, then $y$ is set to $0.5$. As feedback is collected, it is stored as tuples ($\sigma^{0},\sigma^{1},y$) in the reward model data set $D_{RM}$. In general, if $\sigma^{i}>\sigma^{j}$, then the segment $\sigma^{i}$ is preferred by the teacher over segment $\sigma^{j}$. We follow the Bradley-Terry model (Bradley & Terry, 1952) to define a preference predictor using the reward function $\hat{r}_{\theta}$:
@@
Intuitively, this model assumes that the probability of the teacher preferring a segment depends exponentially on the total sum of predicted rewards along the segment. To train the reward function, we can use supervised learning where the teacher provides the labels $y$. More specifically, we update $\hat{r}_{\theta}$ by minimizing the standard binary cross-entropy objective:
$L^{CE}(\theta,D)=-\ E_{(\sigma^{0},\sigma^{1},y)\sim D}\bigl{[}(1-y)\>\text{log}P_{\theta}(\sigma^{0}>\sigma^{1})+\;y\>\text{log}P_{\theta}(\sigma^{1}>\sigma^{0})\bigr{]}$ (2)
@@
The primary difference between scalar and preference-based reward learning is that in scalar-based learning, the human teacher assigns numerical ratings to trajectory segments. In this setting, the comparisons between segments are implicit. More concretely, a teacher assigns a scalar value $y$ to a segment $\sigma^{i}$, and as feedback is collected, it is stored as tuples ($\sigma^{i},y$) in the reward model data set $D_{RM}$. We then apply standard regression and update $\hat{r}_{\theta}$ by minimizing the mean squared error:
$L^{MSE}(\theta,D)=E_{(\sigma^{i},y)\sim D}\bigl{[}(y-\sum_{t}\hat{r}_{\theta}(s^{i}_{t},a^{i}_{t}))^{2}\bigr{]}$ (3)
Human Studies in Human in the Loop RL
To evaluate human-in-the-loop RL algorithms, a common protocol is the use of simulated teachers, where feedback is provided according to a ground truth reward function. This can be useful, as it offers an efficient and controlled evaluation setting. However, we argue that it is essential to evaluate human-in-the-loop RL algorithms with real human feedback. This is especially important in light of recent work finding discrepancies between preference learning algorithms when using simulated teacher feedback versus human feedback (Metcalf et al., 2024).
@@ 4 Sub-optimal Data Pre-training
In this section, we present SDP, a tool that leverages sub-optimal trajectories to improve the feedback efficiency for human-in-the-loop RL. We refer to sub-optimal trajectories as sequences of $(s,a)$ pairs such that:
$r(s_{t},a_{t})-r_{\text{min}}<\epsilon\quad\forall t\in[t,t+k]\quad\text{for some small}\quad\epsilon>0,$ (4)
where $k$ is the segment size and $r_{\text{min}}$ is the minimum possible environment reward. Equation 4 essentially expresses that the rewards achieved along a sub-optimal trajectory should be close to $r_{\text{min}}$. However, it is important to note that in practice we do not have access to the ground truth reward. This prevents us from directly identifying sub-optimal trajectories using Equation 4. Instead, we rely on selecting trajectories that we estimate will align with this criterion, such as gathering trajectories via a random policy.
@@
Once sub-optimal trajectories are collected, we take the approach of pseudo-labeling all transitions with $r_{\text{min}}$. The goal of SDP is then to use this pseudo-labeled data to create a prior for reward models in human-in-the-loop RL methods (see Figure 1). Its simplicity enables SDP to be used in conjunction with any off-the-shelf human-in-the-loop RL algorithm that learns a reward function from human feedback.
@@
SDP comprises two phases: (1) the reward model pre-training phase and (2) the agent update phase. In the reward model pre-training phase, we first gather a data set, $D_{\text{sub}}$, of $N$ sub-optimal state, action transitions. We then pseudo-label all transitions in $D_{\text{sub}}$ with rewards of $r_{\text{min}}$, resulting in $D_{\text{sub}}=\{s_{i},a_{i},r_{\text{min}}\}^{N}_{i=1}$. $D_{\text{sub}}$ is then used to optimize the reward model $\hat{r}_{\theta}$ with the mean squared loss in Equation 3. As a result, the reward model $\hat{r}_{\theta}$ learns to associate all sub-optimal transitions with a low reward. Without such a prior, the reward model would initially have random estimates for these transitions; while the only way to improve such estimates is to obtain feedback from a teacher. Therefore, the reward model pre-training phase provides a valuable reward initialization without requiring any human feedback.
@@
Next, in the agent update phase, we initialize the RL agent’s replay buffer $D_{\text{agent}}$ with $D_{\text{sub}}$. The RL agent then briefly interacts with its environment and performs gradient updates according to its loss functions. The agent update process changes the RL agent’s policy and generates new transitions, which are then stored in both the agent’s replay buffer $D_{\text{agent}}$ and the reward model’s data set $D_{\text{RM}}$. It is important to note that in standard scalar- and preference-based reward learning, we query the teacher for feedback on trajectory segments sampled from $D_{\text{RM}}$. Therefore, adding new transitions into $D_{\text{RM}}$ during the agent update phase is necessary to ensure that the teacher does not provide redundant feedback to the original sub-optimal transitions (as $D_{\text{RM}}$ was empty prior to the agent update phase). When it is time for the teacher to provide their first set of feedback, the feedback can cover a different region of the state and action space, relative to the original sub-optimal data. In Appendix D.1, Figure 8 we empirically show that the agent update phase changes the RL agent’s policy by performing policy rollouts and analyzing the differences in state distributions. See Algorithm 1 for the complete pseudocode.
@@
At first glance, labeling sub-optimal transitions with an incorrect reward may seem problematic, as incorrect labels do introduce statistical bias into the reward model and the RL agent’s value network. However, as the transitions are sub-optimal, we observe that the bias for using an incorrect reward is low (see Figure 9 in Appendix D.2). Moreover, by using the sub-optimal transitions, we increase the overall amount of data used by both models, which can decrease the models’ variance, as shown in the offline RL setting (Yu et al., 2022).
@@
Figure 1: Overview of SDP: After obtaining a data set of sub-optimal trajectories, we pseudo-label the transitions with rewards of $r_{\text{min}}$ (e.g., $r_{\text{min}}=0$). We then pre-train the reward model $\hat{r}_{\theta}$ using this data set. During the agent update phase, we initialize the RL agent’s replay buffer with the same pseudo-labeled data set. The agent then interacts in the environment and makes learning updates to obtain new behaviors for a teacher to give feedback.
@@
Require: Reward model $\hat{r}_{\theta}\leftarrow\theta$ randomly initialized, Reward model data set $D_{\text{RM}}\leftarrow\emptyset$, RL agent with replay buffer $D_{\text{agent}}\leftarrow\emptyset$, Sub-optimal data set $D_{\text{sub}}$ with reward labels $r_{\text{min}}$
@@
3:           Optimize $\hat{r}_{\theta}$ on $D_{\text{sub}}$ with $L^{MSE}$ (Equation 3)
@@
13:  Begin scalar- or preference-based RL using pre-trained $\hat{r}_{\theta}$, RL agent, and $D_{\text{agent}}$, $D_{\text{RM}}$